# The Superclouds of the Local Milky Way

### v1

@@ 6 Conclusion
The most significant outcome of this study is the discovery of a remarkably organized gas distribution within the local kiloparsec of the Milky Way. This distribution follows a distinct pattern of quasi-parallel gas lanes extending over several kiloparsec (see right panel of Fig. 4). These lanes contain substantial masses of the order of 106 M, and exhibit pitch angles of approximately 30. Given their size, mass, and likely composition (predominantly atomic hydrogen), we refer to these structures as superclouds, following the terminology of Elmegreen & Elmegreen (1983). This identification has only become possible with the advent of large-scale 3D dust maps based on data from the Gaia mission (Gaia Collaboration et al., 2023), allowing us to trace the local gas distribution in three dimensions with unprecedented detail. Unlike earlier studies of superclouds relying on Position-Position-Velocity analysis, these new maps offer a clearer, spatially resolved view of the interstellar medium.
The discovery of a coherent network of superclouds within the local kiloparsec marks a step forward in our understanding of the Milky Way’s ISM. These massive, elongated structures appear to serve as scaffolds for molecular cloud formation and star formation on Galactic scales. Much remains to be understood: what these structures truly are, how they form, and how they give rise to GMCs and, ultimately, stars. Their morphology, stability, and apparent independence from the traditional spiral arm framework raise important questions about the large-scale dynamics of the Galaxy. As new data and analysis tools continue to improve our view of the three-dimensional ISM, these findings offer a foundation for future investigations into the interplay between galactic structure, gas dynamics, and star formation. As we refine our three-dimensional view of the nearby Galaxy, we are beginning to define a solid benchmark for connecting the complex structure of the Milky Way with the broader context of galaxy evolution and star formation in the universe.
@@ Acknowledgements.
The authors thank Tine Colman for helpful comments on the HOP algorithm, and Alvaro Hacar for insightful discussions on the topic of superclouds. Co-funded by the European Union (ERC, ISM-FLOW, 101055318). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The results in this paper were based on observations obtained with Gaia, ESA science missions with instruments and contributions directly funded by ESA Member States, NASA, and Canada. This research has made use of NASA’s Astrophysics Data System, the SIMBAD database, and the VizieR catalog access tool operated at CDS, Strasbourg, France, as well as publicly available software libraries, including Astropy (Astropy Collaboration et al., 2013), NumPy (Harris et al., 2020), scikit-learn (Pedregosa et al., 2011), matplotlib (Hunter, 2007), and plotly (Inc., 2015). The computational results have been achieved using the Austrian Scientific Computing (ASC) infrastructure.
@@ A.1 Scikit-Image
Scikit-Image (van der Walt et al. 2014) is a tool for processing and analyzing 3D data sets, making it possible to extract higher density features in the dust data and labeling connected regions. It offers a fast and straightforward segmentation approach, but has a few limitations that make it unsuitable for this work. First, the method requires the minimum density threshold, the only parameter the segmentation depends on, to be predefined. The data is then segmented by distinguishing between high- and low-density voxels, depending on the given threshold. Although the labeling option makes it possible to identify connected structures, the approach is too simple to allow for any fine-tuning.
@@ A.2 Dendrograms
Astrodendro (Robitaille et al. 2019) is a package for constructing dendrograms; hierarchical structures used to identify and analyze nested features. While dendrograms are effective for analyzing and identifying hierarchical substructures, the size of the dust map makes it difficult to analyze regions of interest, due to the large number of resulting branches. This can be avoided by fine-tuning the parameters, but the more effective way is to extract smaller regions for the analysis. However, turning the map into smaller segments is not the preferred approach.
@@ A.3 Convolutional Neural Network
The convolutional neural network (CNN) approach is based on Kim et al. (2020), who developed the algorithm for unsupervised image segmentation in 2D. It was later modified at the DS@univie (private communication with Amanda de Paula Cristino Hirschl) for 3D applications. Originally adapted to find gaps in the dust map by Leike et al. (2020), it can also be applied to detect high-density regions. In this approach, the CNN processes the dust data by extracting structural features based on density. This is done through convolutional layers, which enable the network to learn spatial patterns on different scales. The output is a binary dataset that indicates whether a data point belongs to a structure or is rejected as part of the background. The CNN enables a data-driven segmentation that does not depend on an external density threshold definition, but is not suitable to further segment the defined large-scale structures. This limitation makes it difficult to properly segment the dust map.
@@ Appendix B Cloud Numbering
Figure 7: Same as Fig. 1 (left), with additional cloud ID numbers.
Figure 8: Same as Fig. 4 (left), with additional supercloud ID numbers. ID 0 is the Racliffe wave, 1 th
Table 1: HOP cloud components for each supercloud. The ID refers to the cloud ID as shown in Fig. 7.
@@
0 Radcliffe Wave 0, 12, 18, 27, 30, 34, 37
1 Split 3, 7, 8
2 Natrix Cloud 6, 14, 22, 28, 31, 32, 35
3 Malpolon Cloud 9, 10, 13, 15, 16, 17, 20
25, 19, 24, 36, 38
4 Vela Ridge Cloud 2, 21, 23, 26, 29, 39
@@ Appendix C Cloud Properties
Multiple properties are calculated for the HOP clouds as described in Sect. 3 and can be found in Table 2. The supercloud properties are shown in Table 3. The resulting fit parameters for the Malpolon cloud, Natrix cloud, Radcliffe wave and the Vela Ridge cloud are shown in Table 4.
Table 2: Selection of parameters calculated for the 40 HOP clouds. Each cloud is assigned a cloud ID and has listed which supercloud it is part of. The column cell count indicates the voxels per structure, length (L) gives the cloud length calculated with PCA. The columns X, Y, and Z are the unweighted mean coordinates per cloud. Alpha $\alpha$ and beta $\beta$ represent the face-on and edge-on angles of the clouds.
@@
Table 3: Selection of parameters calculated for the seven superclouds. Each supercloud is assigned a supercloud ID and name. The column cell count indicates the voxels per structure, length (L) gives the cloud length calculated with PCA. The columns X, Y, and Z are the unweighted mean coordinates per cloud. Alpha $\alpha$ and beta $\beta$ represent the face-on and edge-on angles of the clouds.
@@
Table 4: Fit parameters for the undulating superclouds. The column $a$ denotes the amplitude, $\lambda$ the wavelength, $\phi$ the phase shift, and $d$ the vertical offset. The exponential damping parameters $\epsilon_{a}$ for the amplitude and $\epsilon_{\omega}$ for the frequency are only listed for clouds fitted with a damped sinusoidal, as described in Sect. 3.2.4, instead of a standard sinusoidal function.
@@
   Malpolon Cloud 31.97 1137.40 -3.10 58.15 $-1.18\times 10^{-3}$ $-7.52\times 10^{-5}$
   Natrix Cloud 70.53 1749.47 -0.58 -1.17 $1.59\times 10^{-3}$ $-1.12\times 10^{-3}$
   Radcliffe Wave -87.95 1748.00 -3.09 -13.01 $-1.10\times 10^{-3}$ $2.88\times 10^{-4}$

### v3

@@ 6 Conclusion
The most significant outcome of this study is the discovery of a remarkably organized gas distribution within the local kiloparsec of the Milky Way. This distribution follows a distinct pattern of quasi-parallel gas lanes extending over several kiloparsec (see the right panel of Fig. 4). These lanes contain substantial masses, on the order of 106 M, and exhibit pitch angles of approximately 30. Given their size, mass, and likely composition (predominantly atomic hydrogen), we refer to these structures as superclouds, following the terminology of elmegreen1983. This identification has only become possible with the advent of large-scale 3D dust maps based on data from the Gaia mission (gaiaDR3), which allow us to trace the local gas distribution in 3D with unprecedented detail. Unlike earlier studies of superclouds that relied on position-position-velocity analysis, these new maps offer a clearer, spatially resolved view of the ISM.
The discovery of a coherent network of superclouds within the local kiloparsec marks a step forward in our understanding of the Milky Way’s ISM. These massive, elongated structures appear to serve as scaffolds for molecular cloud formation and star formation on Galactic scales. Much remains to be understood: what these structures truly are, how they form, and how they give rise to GMCs and, ultimately, stars. Their morphology, stability, and apparent independence from the traditional spiral arm framework raise important questions about the large-scale dynamics of the Galaxy. As new data and analysis tools continue to improve our view of the 3D ISM, these findings offer a foundation for future investigations into the interplay between galactic structure, gas dynamics, and star formation. As we refine our 3D view of the nearby Galaxy, we are beginning to define a solid benchmark for connecting the complex structure of the Milky Way with the broader context of galaxy evolution and star formation in the Universe.
@@ Acknowledgements.
The authors are particularly grateful to Juan Soler for exciting discussions on the superclouds, for suggesting the HOP algorithm, and for providing helpful comments on the discussion section. We thank Bruce and Debra Elmegreen for insightful discussions on the nature of the superclouds. We also thank Tine Colman for helpful comments on the HOP algorithm, and Alvaro Hacar, Robert Benjamin, and Alyssa Goodman for valuable feedback on the physical state of superclouds. L.A.K. is grateful to Moa Huppenkothen for helpful discussions on the naming of the superclouds. Co-funded by the European Union (ERC, ISM-FLOW, 101055318). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The results in this paper were based on observations obtained with Gaia, an ESA science mission with instruments and contributions directly funded by ESA Member States, NASA, and Canada. This research has made use of NASA’s Astrophysics Data System, as well as publicly available software libraries, including Astropy (robitaille_astropy_2013), NumPy (numpy), scikit-learn (scikit-learn), matplotlib (matplotlib), and plotly (plotly). The computational results have been achieved using the Austrian Scientific Computing (ASC) infrastructure.
@@ A.1 Scikit-Image
Scikit-Image (scikitimage) is a tool for processing and analyzing 3D datasets, making it possible to extract higher density features in the dust data and labeling connected regions. It offers a fast and straightforward segmentation approach, but has a few limitations that make it unsuitable for this work. First, the method requires the minimum density threshold, the only parameter the segmentation depends on, to be predefined. The data are then segmented by distinguishing between high- and low-density voxels, depending on the given threshold. Although the labeling option makes it possible to identify connected structures, the approach is too simple to allow for any fine-tuning.
@@ A.2 Dendrograms
Astrodendro (astrodendro) is a package for constructing dendrograms; hierarchical structures used to identify and analyze nested features. While dendrograms are effective for analyzing and identifying hierarchical substructures, the size of the dust map makes it difficult to analyze regions of interest, due to the large number of resulting branches. This can be avoided by fine-tuning the parameters, but the more effective way is to extract smaller regions for the analysis. However, turning the map into smaller segments is not the preferred approach.
@@ A.3 Convolutional neural network
The convolutional neural network (CNN) approach is based on kim2020, who developed the algorithm for unsupervised image segmentation in 2D. It was later modified at the DS/Univie (private communication with Amanda de Paula Cristino Hirschl) for 3D applications. Originally adapted to find gaps in the dust map by Leike2020, it can also be applied to detect high-density regions. In this approach, the CNN processes the dust data by extracting structural features based on density. This is done through convolutional layers, which enable the network to learn spatial patterns on different scales. The output is a binary dataset that indicates whether a data point belongs to a structure or is rejected as part of the background. The CNN enables a data-driven segmentation that does not depend on an external density threshold definition, but is not suitable to further segment the defined large-scale structures. This limitation makes it difficult to properly segment the dust map.
@@ Appendix B Cloud numbering
Figure 7: Same as Fig. 1 (left) but with additional cloud ID numbers.
Figure 8: Same as Fig. 4 (left) but with additional supercloud ID numbers.
Table 1: HOP cloud components for each supercloud; the ID refers to the cloud ID as shown in Fig. 7.
@@
0 Radcliffe Wave 0, 37, 18, 12, 30, 27, 34
1 Split 8, 3, 7
2 Natrix Cloud 22, 6, 14, 31, 28, 35, 32
3 Malpolon Cloud 24, 10, 25, 9, 17, 36, 16
13, 15, 20, 38, 19
4 Vela Ridge Cloud 26, 23, 29, 39, 2, 21
@@ Appendix C Cloud properties
Multiple properties are calculated for the HOP clouds as described in Sect. 3 and can be found in Table 2. The supercloud properties are shown in Table 3. The resulting fit parameters for the Malpolon Cloud, Natrix Cloud, Radcliffe Wave, and the Vela Ridge Cloud are shown in Table 4.
Table 2: Selection of parameters calculated for the 40 HOP clouds. [^2] Each cloud is assigned a cloud ID and has listed which supercloud it is part of. The column cell count indicates the voxels per structure, L gives the cloud length calculated with PCA. The columns X, Y, and Z are the unweighted mean coordinates per cloud. The angles $\alpha$ and $\beta$ represent the face-on and edge-on angles of the clouds.
@@
Table 3: Selection of parameters calculated for the seven superclouds. [^3] Each supercloud is assigned a name. The column cell count indicates the voxels per structure, L gives the cloud length calculated with PCA. The columns X, Y, and Z are the unweighted mean coordinates per cloud. The angles $\alpha$ and $\beta$ represent the face-on and edge-on angles of the clouds.
@@
Table 4: Fit parameters for the undulating superclouds.  [^4] The column $a$ denotes the amplitude, $\lambda$ the wavelength, $\phi$ the phase shift, and $d$ the vertical offset. The exponential damping parameters $\epsilon_{a}$ for the amplitude and $\epsilon_{\omega}$ for the frequency are only listed for clouds fitted with a damped sinusoidal, as described in Sect. 3.2.4, instead of a standard sinusoidal function.
@@
  Malpolon Cloud 31.97 1137.40 -3.10 58.15 $1.18\times 10^{-3}$ $7.52\times 10^{-5}$
  Natrix Cloud 70.53 1749.47 -0.58 -1.17 $59\times 10^{-3} 1.1$$12\times 10^{-3}$
  Radcliffe Wave -87.95 1748.00 -3.09 -13.01 $1.10\times 10^{-3}$ $2.88\times 10^{-4}$
@@
Appendix D Smaller-scale clouds
The final segmentation parameters were selected to identify large-scale structures within the dust map. Consequently, several smaller clouds that did not satisfy the final HOP criteria for minimum size or density threshold were excluded. While these clouds are not the focus of this study, they highlight the importance of studying smaller-scale structures alongside the superclouds. Complementary work on smaller-scale structures can be found in xie2024. To focus on smaller-scale structures, we also performed several test runs with different parameter sets, revealing interesting features absent from the final segmentation.
One of the most prominent features when using a lower mincells value of 30.000 is a connection between the Malpolon Cloud (HOP cloud 15) and Natrix Cloud (HOP cloud 31). This roughly 200 pc long connection is visible by eye in the Edenhofer2024 dust map at approximately X = -730 pc and Y = 60 pc. The origin of this feature is currently unknown, a possible explanation is the formation through external influences such as feedback processes, or it may indicate that the two superclouds were more strongly connected in the past. Further investigation is needed to understand the origin and significance of this feature.
Another cloud recovered in this segmentation is located below HOP cloud 20 at approximately X = -920 pc and Y = 120 pc. While this cloud does not appear to be associated with any of the superclouds, it is a prominent feature in the Edenhofer2024 dust map. Given its compact and isolated appearance, it may be of interest for studies focused on smaller-scale structures of higher density.
With the lower mincells value, we also recover the top of the Pegasus Ridge (Swiggum et al. in prep). This cloud lies below the Radcliffe Wave, with the top part appearing to wrap around the edge of the Local Bubble (oneill2024). As shown in Swiggum et al. (in prep), it further extends below the Radcliffe Wave, but the lower part is composed of material with densities too low to be recovered by HOP.
Reducing the mincells parameter further leads to the recovery of additional small-scale clouds. However, these features do not affect the overall large-scale supercloud structure and lie outside the scope of the present study.
Appendix E Superclouds without sinusoidal fits
The three remaining superclouds, which are not shown with a sinusoidal fit in Fig. 5, are shown in Fig. 9. As mentioned, the SSE and the Anguis Cloud are excluded from the analysis, as the vertical displacement of the clouds is not as prominent as in the other examples. The Split is excluded, as it appears unusually flat, with only two noticeable downward extensions at Y = -200 pc (referred to as the ”C” and discussed in edenhofer2024_c) and Y = 500 pc.
Figure 9: 2D projection in the Z/Y plane of the three superclouds without a sinusoidal fit.